# DiT-HC: Enabling Efficient Training of Visual Generation Model DiT on HPC-oriented CPU Cluster

### v1

@@ 1. Introduction
Artificial intelligence is becoming increasingly prevalent in scientific research, and scientific computing is shifting from a paradigm centered solely on numerical simulation to one where simulations and data-driven models co-develop in a complementary manner(silvano2025survey; ejarque2022enabling; jiang2024efficient). While this shift enhances modeling capability(bracco2025machine) and improves the efficiency of data utilization, it also places new demands on high-performance computing (HPC) platforms: they must not only continue to support large-scale simulations in domains such as climate and seismology, but also run generative foundation models efficiently to enable data-driven optimization in areas such as data assimilation(huang2024diffda; xu2025fuxi), parameterization schemes(chen2023resu; duan2025ai), and bias correction(kim2021deep; gregory2024machine). Consequently, HPC platforms are supposed to support both numerical simulation and model training to achieve a unified workflow.
@@
HPC-oriented CPU cluster is one of the most established and widely deployed computing infrastructures in modern supercomputers, forming the backbone of large-scale numerical simulations and scientific applications. In generative model training, GPUs/TPUs are the primary platforms(li2022ai), while CPU clusters provide only limited support for efficient training of deep learning models. The primary reason for this gap lies in the general-purpose nature of CPUs: their architectures emphasize versatility but typically provide lower compute density and memory bandwidth than GPUs. For deep learning workloads dominated by large-scale matrix operations and intensive memory access workloads, these limitations directly constrain training efficiency, placing CPUs at a disadvantage in scaling generative models and posing challenges unifying AI and HPC on the same platform.
@@
Consequently, there is a trend for next-generation HPC-oriented CPU architectures to incorporate AI-focused hardware features, such as matrix acceleration units (MAUs) and on-package high-bandwidth memory (OPM), which provide new opportunities for deep learning model training. For example, the Armv9 architecture introduces two-dimensional ZA registers and the Scalable Matrix Extension (SME) instruction set, significantly improving the efficiency of matrix and tensor computations (weidmann2021introducing). Intel’s Sapphire Rapids processors similarly integrate Advanced Matrix Extensions (AMX), enabling more efficient execution of AI workloads (kim2024exploiting). In addition, modern HPC CPUs are adopting hierarchical memory structures that combine on-package high-bandwidth memory with traditional DDR, alleviating the bandwidth bottlenecks of conventional designs. However, fully exploiting these new features is non-trivial: unlike the highly encapsulated GPU ecosystem with its mature toolchains, CPU platforms involve distinct characteristics in instruction set extensions, hierarchical memory management, and NUMA-domain interconnects, which necessitate systematic rethinking of operator implementations, parallelization strategies, and memory scheduling.
@@
In this context, we select the generative foundation model DiT as our case study to investigate efficient training and scalable optimization on HPC-oriented CPU clusters. On the one hand, diffusion models have become the dominant paradigm in visual content generation (sohl2015deep; ho2020denoising; song2020score; nichol2021improved; lu2022dpm; podell2023sdxl; rombach2022text), and Transformer-based DiT (Peebles2022DiT; esser2024scaling) serves as a representative with strong global modeling capacity and scalability. On the other hand, diffusion models have also demonstrated significant value in scientific domains (kazerouni2023diffusion), achieving advances in satellite image super-resolution (dong2024building), multimodal image fusion (bao2023one), ensemble weather forecasting (price2023gencast), and Earth system data assimilation (hess2025fast; huang2024diffda; springenberg2025diffscale). Therefore, DiT combines the representativeness of cutting-edge generative models with practical relevance in scientific applications, making it a representative target for evaluating and optimizing next-generation CPUs in large-scale training.
@@ 2. Related Works
Several attempts have explored deep learning training on CPU-based supercomputers. The Fugaku system (r1_dongarra2020report) with A64FX processors optimized oneDNN to scale applications (r2_sato2020co), while the Sunway Taihulight system (fu2016sunway) employed swCaffe (li2018swcaffe) for distributed training. CPUs have also been used for memory-intensive tasks: Huang (huang2021hierarchical) proposed hierarchical training for recommendation models, and Intel researchers (kalamkar2020optimizing) optimized deep recommendation workloads on Skylake/Cascade Lake CPUs. At large scale, You et al. (you2018imagenet) trained AlexNet in 11 minutes using 2,048 Xeon CPUs with LARS. These works demonstrate the feasibility of CPU-based training, particularly for memory-demanding workloads, but they mainly targeted non-Transformer models that are less representative in today’s large-model landscape, and their optimizations were highly case-specific rather than generalizable.
At the operator level, high-performance math libraries provide essential support. oneDNN (r3_onednn) accelerates deep learning across multiple platforms, while Arm-specific efforts such as autoGEMM (r5_wu2024autogemm), LBBGEMM (r6_wei2022lbbgemm), and automatic micro-kernel generation (r4_alaejos2024automatic) optimize GEMM computations through dynamic tiling, auto-tuning, or code generation. Although these methods achieve strong performance on traditional CPUs, they remain focused on vector-based instruction sets and fail to exploit newer architectural features such as matrix acceleration units (MAUs) and on-package high-bandwidth memory, limiting their effectiveness on processors like LX2.
At the framework level, diffusion and Transformer models have motivated specialized optimizations. DiffusionPipe (r10_tian2024diffusionpipe) explored pipeline-parallel training strategies for DiT, while general Transformer frameworks such as xFormers (r8_xFormers2022), MEATTEN (r11_fu2024optimizing), OneFlow (r9_yuan2021oneflow), DeepSpeed (rasley2020deepspeed), and Megatron (shoeybi2019megatron) introduced operator fusion, modular attention kernels, and distributed scaling methods. Among them, MEATTEN targets Arm CPUs with fusion and tiling optimizations. However, the majority of these frameworks are GPU-centric, making them difficult to port to CPU clusters. This gap highlights the need for CPU-specific frameworks that can exploit new architectural features while supporting representative Transformer-based generative models such as DiT.

### v2

@@ 1. Introduction
Artificial intelligence is becoming increasingly prevalent in scientific research, and scientific computing is shifting from a paradigm centered solely on numerical simulation to one where simulations and data-driven models co-develop in a complementary manner(Silvano et al., 2025; Ejarque et al., 2022; Jiang et al., 2024). While this shift enhances modeling capability(Bracco et al., 2025) and improves the efficiency of data utilization, it also places new demands on high-performance computing (HPC) platforms: they must not only continue to support large-scale simulations in domains such as climate and seismology, but also run generative foundation models efficiently to enable data-driven optimization in areas such as data assimilation(Huang et al., 2024; Xu et al., 2025), parameterization schemes(Chen et al., 2023; Duan et al., 2025), and bias correction(Kim et al., 2021; Gregory et al., 2024). Consequently, HPC platforms are supposed to support both numerical simulation and model training to achieve a unified workflow.
@@
HPC-oriented CPU cluster is one of the most established and widely deployed computing infrastructures in modern supercomputers, forming the backbone of large-scale numerical simulations and scientific applications. In generative model training, GPUs/TPUs are the primary platforms(Li et al., 2022), while CPU clusters provide only limited support for efficient training of deep learning models. The primary reason for this gap lies in the general-purpose nature of CPUs: their architectures emphasize versatility but typically provide lower compute density and memory bandwidth than GPUs. For deep learning workloads dominated by large-scale matrix operations and intensive memory access workloads, these limitations directly constrain training efficiency, placing CPUs at a disadvantage in scaling generative models and posing challenges unifying AI and HPC on the same platform.
@@
Consequently, there is a trend for next-generation HPC-oriented CPU architectures to incorporate AI-focused hardware features, such as matrix acceleration units (MAUs) and on-package high-bandwidth memory (OPM), which provide new opportunities for deep learning model training. For example, the Armv9 architecture introduces two-dimensional ZA registers and the Scalable Matrix Extension (SME) instruction set, significantly improving the efficiency of matrix and tensor computations (Weidmann, 2021). Intel’s Sapphire Rapids processors similarly integrate Advanced Matrix Extensions (AMX), enabling more efficient execution of AI workloads (Kim et al., 2024). In addition, modern HPC CPUs are adopting hierarchical memory structures that combine on-package high-bandwidth memory with traditional DDR, alleviating the bandwidth bottlenecks of conventional designs. However, fully exploiting these new features is non-trivial: unlike the highly encapsulated GPU ecosystem with its mature toolchains, CPU platforms involve distinct characteristics in instruction set extensions, hierarchical memory management, and NUMA-domain interconnects, which necessitate systematic rethinking of operator implementations, parallelization strategies, and memory scheduling.
@@
In this context, we select the generative foundation model DiT as our case study to investigate efficient training and scalable optimization on HPC-oriented CPU clusters. On the one hand, diffusion models have become the dominant paradigm in visual content generation (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Nichol and Dhariwal, 2021; Lu et al., 2022; Podell et al., 2023; Rombach et al., 2022), and Transformer-based DiT (Peebles and Xie, 2022; Esser et al., 2024) serves as a representative with strong global modeling capacity and scalability. On the other hand, diffusion models have also demonstrated significant value in scientific domains (Kazerouni et al., 2023), achieving advances in satellite image super-resolution (Dong et al., 2024), multimodal image fusion (Bao et al., 2023), ensemble weather forecasting (Price et al., 2023), and Earth system data assimilation (Hess et al., 2025; Huang et al., 2024; Springenberg et al., 2025). Therefore, DiT combines the representativeness of cutting-edge generative models with practical relevance in scientific applications, making it a representative target for evaluating and optimizing next-generation CPUs in large-scale training.
@@ 2. Related Works
Several attempts have explored deep learning training on CPU-based supercomputers. The Fugaku system (Dongarra, 2020) with A64FX processors optimized oneDNN to scale applications (Sato et al., 2020), while the Sunway Taihulight system (Fu et al., 2016) employed swCaffe (Li et al., 2018) for distributed training. CPUs have also been used for memory-intensive tasks: Huang (Huang et al., 2021) proposed hierarchical training for recommendation models, and Intel researchers (Kalamkar et al., 2020) optimized deep recommendation workloads on Skylake/Cascade Lake CPUs. At large scale, You et al. (You et al., 2018) trained AlexNet in 11 minutes using 2,048 Xeon CPUs with LARS. These works demonstrate the feasibility of CPU-based training, particularly for memory-demanding workloads, but they mainly targeted non-Transformer models that are less representative in today’s large-model landscape, and their optimizations were highly case-specific rather than generalizable.
At the operator level, high-performance math libraries provide essential support. oneDNN (27) accelerates deep learning across multiple platforms, while Arm-specific efforts such as autoGEMM (Wu et al., 2024), LBBGEMM (Wei et al., 2022), and automatic micro-kernel generation (Alaejos et al., 2024) optimize GEMM computations through dynamic tiling, auto-tuning, or code generation. Although these methods achieve strong performance on traditional CPUs, they remain focused on vector-based instruction sets and fail to exploit newer architectural features such as matrix acceleration units (MAUs) and on-package high-bandwidth memory, limiting their effectiveness on processors like LX2.
At the framework level, diffusion and Transformer models have motivated specialized optimizations. DiffusionPipe (Tian et al., 2024) explored pipeline-parallel training strategies for DiT, while general Transformer frameworks such as xFormers (Lefaudeux et al., 2022), MEATTEN (Fu et al., 2024), OneFlow (Yuan et al., 2021), DeepSpeed (Rasley et al., 2020), and Megatron (Shoeybi et al., 2019) introduced operator fusion, modular attention kernels, and distributed scaling methods. Among them, MEATTEN targets Arm CPUs with fusion and tiling optimizations. However, the majority of these frameworks are GPU-centric, making them difficult to port to CPU clusters. This gap highlights the need for CPU-specific frameworks that can exploit new architectural features while supporting representative Transformer-based generative models such as DiT.